UT 20130131 Technical Observing

Observer: JHill (Tucson)
Telescope Operator: DHuerta (LBTO)
Additional AO Support: DMiller (LBTO)
Telescope Support: JLittle (LBTO)
Instrument Support: JMorris (LBTO)


This night was lost to weather due to significant rime ice remaining on the dome. A portion of the night was given over to closed dome testing of Build 2013A.


Startup Issues

Authorized at LUCI1 @ LFBG......

01:30 AGW1 AzCams are now working fine after a power cycle and reboot of the computers.

02:00 Fixed up some details of AGwStages and AzCamServer pages on the wiki.

02:10 LFBG rotator giving us some issues - showing e-stop.

Closed Dome Presets


02:56 ACQUIRE preset

02:58 TRACK preset

Turn on fake guide stars, restart GCSL and GCSR.

03:03 ACQUIRE preset

03:05 GUIDE preset

03:08 offset 10 0 (or similar) Probe moved -0.0685539, 28.616 to -7.60586, 32.4467.

03:10 ACTIVE preset


03:12 ACQUIRE preset

03:13 GUIDE preset

03:15 ACTIVE preset

03:15 offset 10 0

MODS1 AGw is turned off.


Found IRTC cooling TEC had been left on.

Not communicating correctly with AGW2. Had to restart GCSR.

AGW2 has "classical" network error. Restart oacserver. Restart GCSR again.

03:23 takepic 0.1 with IRTC.

03:29 Preset fails because DX M3 brake not set. Something has caused it to lose the home knowledge. Was there a power outage?


03:32 LBC BOTH TRACK preset to WT10_368.

03:35 DOFPIA sends wavefront command to both sides.

Turned fake guide stars off. Restart GCSs.

Turned LBC off.

03:40 Park the telescope.

-- JohnHill - 31 Jan 2013
Topic revision: r3 - 05 Feb 2013, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback